OpenStudy (anonymous):

At a certain factory 5/8 of the workers are married and 3/4 are at least 40 years old. If 1/4 of the married workers are younger than 40, what fraction of the workers at least 40 are not married?

OpenStudy (anonymous):

lets c 3/8ths of the company isnt married (1/4)(5/8)=5/32 of the company is not married and under 40 1/4 of the whole company is under 40 so 1/4-5/32=amount of ppl under 40 and not married so its 3/32 for sure